0 a o GDP he s CD plyer dses not ceetribute to GD as 3 PROBLEMS AND APPLICATIONS (S0 marks ) hems is worth in terms of todays dollars using 177 as the price index for today. C15 marks) CPi was 17.7 and the price of a movie ticket was $0.25 h. In 1952, the CPi was 13.1 anda cook earmed $15.00 a week c. In 1943, he CPl was 174 and a gallon of gas cost So.19
Answer #1

a.Price of a Movie ticket in terms of today's dollars = 0.25*177/17.7 = $2.5

b. Cook earning a week in terms of today's dollars = 15*177/13.1 = $202.67

c. A gallon of gas cost in terms of today's dollars = 0.19*177/17.4 = $1.93